Fan-Tastic Sports Gear recorded $2,900,000 of sales last year and projects sales to increase by $360,000 in the current year. Last year, 90% of sales were on account, with over 300 customer accounts. Bad debt expense was $26,187.
1. Assume that Fan-Tastic Sports Gear used the allowance method last year, and the allowance account at the end of the year had a debit balance of $2,190. The company estimated uncollectible accounts expense using the percent of credit sales method and expected 0.75% of credit sales to be uncollectible. What is the amount of the adjusting entry to provide for doubtful accounts on December 31? Round all computations to the nearest dollar.
2. How much higher (lower) would Fan-Tastic Sports Gear’s net income have been under the allowance method assumption in (1) than under the direct write-off method? (Enter “0” if there is no change.) by
3. Using the allowance method, the net realizable value of the receivables would appear on which financial statement?
Answer #1

1. Adjusting Entry Amount

Expected sales in Current year (a) = ($ 2,900,000+ $360,000) = $ 3,260,000

Expected Credit sales (b=a*90%) = $ 2,934,000

Expected Uncollectable Debt (c=b*.75%) = $ 22,005

Avaiable balance in Allowence Account (d) = $ 2,190

Adjusting entry amount/Allocation from profit and loss account for the period (c-d) = $19,815

2. 0

The direct write-off method recognizes bad accounts as an expense at the point when judged to be uncollectible and is the required method for federal income tax purposes. The allowance method provides in advance for uncollectible accounts think of as setting aside money in a reserve account.

Hence if direct mathod were followed the written off will happen when it actually incures , in the absense of actual bad debt incurred in the year i assume that there will be $22,005 will be the bad debt so there will be no difference

3.In Balance sheet

Individual balances are generated by sales made on credit. According to U.S. GAAP, the figure that is presented on a balance sheet for accounts receivable is its net realizable value—the amount of cash the company estimates will be collected over time from these accounts